Christine's age is 6 more than ken's age. Three times ken's age plus 5 times Christina's age is equal to 46.how old is christine?
Answer by ikleyn(52803)

.
Christine's age is 6 more than ken's age. Three times ken's age plus 5 times Christina's age is equal to 46.how old is christine?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

C = K + 6,     (1)
3K + 5C = 46.  (2)

Substitute (1) into (2), simplify and solve.

3K + 5*(K+6) = 46,

3K + 5K + 30 = 46,

8K = 46 - 30 = 16,

K = 2, C = K + 6 = 2+6 = 8.

Answer. K =2, C = 8.